# RC-positivity, comparison theorems and prescribed Hermitian-Yang-Mills tensors I

### 11 Mar 2026 v1

@@ 1. Introduction
The method for solving equation (1.4) differs significantly from those for equations (1.1) and (1.2), and we provide a brief overview of the key ideas involved. The uniqueness part of Theorem 1.1 is established by the following comparison theorem for Hermitian-Yang-Mills tensors of Hermitian vector bundles:
@@
Note that the curvature constraint $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$ in Theorem 1.1 is closely related to the notions of RC-positivity and uniform RC-positivity introduced in [Yang18, Yang20]. A comparison theorem analogous to Theorem 1.2 for RC-positive vector bundles was also established in [XYY24+]. This curvature constraint should be replaced by RC-positivity, and this replacement will be carried out in a separate project on general Hermitian manifolds.
@@
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. If there exists a Hermitian metric $h_{0}$ on $E$ such that
@@
(1.12) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
@@ 6. Chern number inequalities
In this section, we prove Theorem 1.6:
@@
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. If there exists a Hermitian metric $h_{0}$ on $E$ such that
@@
(6.2) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
@@
By Theorem 1.1, there exists a smooth Hermitian metric $h$ on $E$ such that
@@
$R^{h}=R^{h}_{i\overline{j}\alpha\overline{\beta}}dz^{i}\wedge d\overline{z}^{i}e^{\alpha}\otimes\overline{e}^{\beta}.$
@@
This completes the proof. ∎
@@ 7. The prescribed Hermitian-Yang-Mills tensor equation on line bundles
$Q(t)$ $=$ $\int_{M}\left(\Delta_{g}f\right)/(e^{f+t}-1)\cdot\omega_{g}^{n}=-\int_{M}g\left(\nabla f,\nabla(e^{f+t}-1)^{-1}\right)\cdot\omega_{g}^{n}$

### 13 Mar 2026 v2

@@ 1. Introduction
By Serre duality, the corresponding conjugate equation of (1.4) admits a solution when $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is (quasi-)negative definite; see Remark 5.2 for further details. The method for solving equation (1.4) differs significantly from those for equations (1.1) and (1.2), and we provide a brief overview of the key ideas involved. The uniqueness part of Theorem 1.1 is established by the following comparison theorem for Hermitian-Yang-Mills tensors of Hermitian vector bundles:
@@
Note that the curvature constraint $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}>0$ in Theorem 1.1 is closely related to the notions of RC-positivity and uniform RC-positivity introduced in [Yang18, Yang20]. A comparison theorem analogous to Theorem 1.2 for RC-positive vector bundles was also established in [XYY24+]. This curvature constraint should be replaced by RC-positivity, and this replacement is carried out in a separate project on general Hermitian manifolds.
@@
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. Suppose that there exists a Hermitian metric $h_{0}$ on $E$ satisfying
@@
for some constants $a,b\in{\mathbb{R}}$. Then the following Chern number inequality holds
(1.12) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n},$
if $ab>0$ or $M$ is projective.
@@ 5. Proofs of main results
Remark 5.2.
The dual equation case can be addressed via Serre duality. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}<0$. Then the dual metric $h_{0}^{*}$ on $E^{*}$ satisfies
$\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}^{*}}>0$
as a Hermitian tensor in $\Gamma(M,E\otimes\overline{E})$ under the natural identification $E\cong(E^{*})^{*}$. Hence, for any Hermitian positive definite tensor $P_{*}\in\Gamma(M,E\otimes\overline{E})$, there exists a unique smooth Hermitian metric $h_{*}$ on $E^{*}$ such that
(5.2) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{*}}=P_{*}.$
Or equivalently, for any Hermitian negative definite tensor $P\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$, there exists a unqiue smooth Hermitian metric $h$ on $E$ satisfying the dual equation:
(5.3) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=H\cdot P\cdot\overline{H^{t}}$
where $H=h\cdot h_{0}^{-1}\in\Gamma(M,E^{*}\otimes E)$.
@@ 6. Chern number inequalities
In this section, we prove Theorem 1.6 in a slightly general setting:
@@
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. Suppose that there exists a Hermitian metric $h_{0}$ on $E$ satisfying
@@
for some constants $a,b\in{\mathbb{R}}$. Then the following Chern number inequality holds
(6.2) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n},$
if $ab>0$ or there exists a holomorphic line bundle $L$ on $M$ satisfying $\int_{M}c_{1}(L)\wedge\omega_{g}^{n}\neq 0$.
@@
We assume that $b\geq a>0$. By Theorem 1.1, there exists a smooth Hermitian metric $h$ on $E$ such that
@@
$R^{h}=R^{h}_{i\overline{j}\alpha\overline{\beta}}dz^{i}\wedge d\overline{z}^{j}\otimes e^{\alpha}\otimes\overline{e}^{\beta}.$
@@
For the case $a\leq b<0$, we know that
(6.12) $0<-b\cdot h_{0}^{*}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}^{*}}\leq-a\cdot h_{0}^{*}$
as Hermitian tensors in $\Gamma(M,E\otimes\overline{E})$ under the isomorphism $E\cong(E^{*})^{*}$. Hence, there exists a Hermitian metric $h^{*}$ on $E^{*}$ such that
(6.13) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h^{*}}=h^{*}_{0}.$
Moreover, one has
(6.14) $0<-b\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h^{*}}\leq-\Lambda_{\omega_{g}}\sqrt{-1}R^{h^{*}_{0}}\leq-a\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h^{*}},$
and so $bh^{*}\leq h_{0}^{*}\leq-ah^{*}$. This implies
(6.15) $\int_{M}\left(2rc_{2}(E^{*})-(r-1)c^{2}_{1}(E^{*})\right)\wedge\omega_{g}^{n-2}\geq-\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
Since $c_{1}^{2}(E^{*})=c_{1}^{2}(E)$ and $c_{2}(E^{*})=c_{2}(E)$, we get the result.
Suppose that there exists a holomorphic line bundle $L$ over $M$ with $\int_{M}c_{1}(L)\wedge\omega_{g}^{n-1}>0$. It is easy to show that there exists a smooth Hermitian metric $h^{L}$ on $L$ such that
(6.16) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h^{L}}=c_{0}=\frac{n\int_{M}c_{1}(L)\wedge\omega_{g}^{n-1}}{\int_{M}\omega_{g}^{n}}>0.$
We consider the twisted Hermitian vector bundle $E_{k}:=E\otimes L^{k}$. The Hermitian-Yang-Mills tensor of $(E_{k},h_{0}\otimes h^{kL})$ is given by
(6.17) $\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}\otimes h^{kL}}\right)=\left(\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\right)\otimes h^{kL}+h_{0}\otimes\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h^{kL}}\right).$
In particular,
(6.18) $\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}\otimes h^{kL}}\right)=\left(\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}+kc_{0}h_{0}\right)\otimes h^{kL}.$
We fix a positive integer $k$ such that $a+kc_{0}>0$. Then the Hermitian-Yang-Mills tensor of $(E_{k},h_{0}\otimes h^{kL})$ satisfies
(6.19) $0<(a+kc_{0})(h_{0}\otimes h^{kL})\leq\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}\otimes h^{kL}}\right)\leq(b+kc_{0})(h_{0}\otimes h^{kL}).$
In particular, by the first case, we deduce that
(6.20) $\int_{M}\left(2rc_{2}(E\otimes L^{k})-(r-1)c^{2}_{1}(E\otimes L^{k})\right)\wedge\omega_{g}^{n-2}\geq-\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
On the other hand, it is well-known that
$\int_{M}\left(2rc_{2}(E\otimes L^{k})-(r-1)c^{2}_{1}(E\otimes L^{k})\right)\wedge\omega_{g}^{n-2}=\int_{M}\left(2rc_{2}(E)-(r-1)c^{2}_{1}(E)\right)\wedge\omega_{g}^{n-2}.$
Hence, we obtain the result. ∎
Remark 6.2.
It is expected that the projectivity condition in Theorem 1.6 is redundant, i.e., the Chern number inequality (1.12) holds for any $a,b\in{\mathbb{R}}$ satisfying (1.11).
@@ 7. The prescribed Hermitian-Yang-Mills tensor equation on line bundles
$2Q(t)$ $=$ $2\int_{M}\left(\Delta_{g}f\right)/(e^{f+t}-1)\cdot\omega_{g}^{n}=-\int_{M}g\left(\nabla f,\nabla(e^{f+t}-1)^{-1}\right)\cdot\omega_{g}^{n}$